# MMLongBench-Doc: Benchmarking Long-context Document Understanding with Visualizations

### v2

@@ Appendix D Limitations
One primary limitation is the scale of the benchmark. Currently, our benchmark includes a test set comprising 130 documents and 1,062 questions. It is much smaller compared to previous datasets. The complexity and difficulty of annotations limit the scale of our benchmark. As a long-context benchmark, our documents average about 50 pages and 20,000 tokens. And most questions require either complicated reasoning or cross-page comprehension. It takes more than one hour for an expert-level annotator to read through a single document, and then edit existing instances and create new instances on this document. Given the purpose of MMLongBench-Doc as an evaluation benchmark, we prioritize annotation quality over quantity. Moreover, the results presented in Sections 3.3 and 3.4 confirm that the scale of our benchmark is sufficient for fine-grained evaluations across different document types, evidence sources, evidence pages, etc.. Additionally, we plan to expand our benchmark by adding more documents and questions in future iterations.

### v3

@@ Appendix D Limitations
One primary limitation is the scale of the benchmark. Currently, our benchmark includes a test set comprising 135 documents and 1,082 questions. It is much smaller compared to previous datasets. The complexity and difficulty of annotations limit the scale of our benchmark. As a long-context benchmark, our documents average about 50 pages and 20,000 tokens. And most questions require either complicated reasoning or cross-page comprehension. It takes more than one hour for an expert-level annotator to read through a single document, and then edit existing instances and create new instances on this document. Given the purpose of MMLongBench-Doc as an evaluation benchmark, we prioritize annotation quality over quantity. Moreover, the results presented in Sections 3.3 and 3.4 confirm that the scale of our benchmark is sufficient for fine-grained evaluations across different document types, evidence sources, evidence pages, etc.. Additionally, we plan to expand our benchmark by adding more documents and questions in future iterations.